Consider a simple paging system with the following parameters: 232 bytes of physical memory; page size...

  1. Consider a simple paging system with the following parameters: 232 bytes of physical memory; page size of 210 bytes; 216 pages of logical address space.
    1. How many entries in the page table?
    1. How many bits in each page table entry? Assume each page table entry contains a valid/invalid bit.
Answer #1

A) physical memory size =2^32 bytes

page size = 2^10 bytes

logical address space : 2^16 bytes

A)

Entries in page table = no of pages in virtual memory

                                = 2^16

==========================================

B)frames in physical memory = size of physical memory/size of the page frame

                                            =2^32/2^10

                                           =2^22

no of bits in physical memory=22

bits in each page table entry when each page table entry contains a valid/invalid bit. =22+1

=23